# StochasticGW-GPU: rapid quasi-particle energies for molecules beyond 10000 atoms

### v1

@@ 1 Introduction
For systems containing thousands of atoms or more, one can evaluate the self-energy operator using a stochastic GW formulation at greatly reduced cost, as detailed by our previous works 48, 49, 50, 51. Here, we briefly summarize the main features of the method. First, we evaluate the self-energy operator in the time domain to exploit the direct product computation of $\varSigma\left(t\right)$ from Green’s function $G$ and screened Coulomb potential $W$; we Fourier transform the resulting $\Sigma\left(t\right)$ to $\Sigma\left(\omega\right)$ only in the final stage of the calculation. Second, we invoke the stochastic Resolution of Identity (sRoI) 52 and define random orbital functions to use as bases for evaluating the Green’s function $G$ and effective polarization $W$. We then compute the expectation values of these operators using real time propagation and accumulate statistical averages over products of random samples. This is the main ingredient of stochastic GW and it has the advantage of allowing one to decouple the spatial- and time- dependence in the six-dimensional integrals needed to evaluate $\Sigma\left(t\right)$ 48. As a result, instead of requiring the full space of occupied and unoccupied orbitals and energies $\left\{\phi_{k},\varepsilon_{k}\right\}$ (which typically number in the tens of thousands or for a thousand-atom molecule), we, in effect, evaluate $G$ and $W$ using compact sets of stochastic linear combinations of the occupied or unoccupied orbital space. An additional benefit of sROI is that computations in the stochastic bases can be done independently, enabling the critical path of the calculation to be made embarrassingly parallelizable. Third, we incorporate sparse stochastic compression 50 in our stochastic Time-Dependent Hartree propagation algorithm 48, 53 for evaluating $W$. This enables computing components of $W$ over a collection of randomly-chosen short segments without needing a full spatial grid, reducing storage costs. Finally, instead of projecting each stochastic sample onto the full set of occupied orbitals $\left\{\phi_{k}^{KS}\right\}_{occ}$ from the preliminary DFT calculation (incurring substantial I/O and computational costs), we filter these samples to generate occupied stochastic orbitals. We construct a filter from a Chebyshev polynomial expansion of the Kohn-Sham Hamiltonian 54. While the filtering approach has the disadvantage of requiring many terms to produce a sharp cutoff at the Fermi energy, we recently found that the expansion length of the filter can be greatly decreased 51 by relaxing the expansion to have zero weights inside the band gap (where no states are present); this, in turn, reduces the number of matrix-vector products needed to prepare the occupied stochastic orbitals.
@@ 2.1 Algorithmic Overview
Our stochastic GW implementation is similar to that described previously 50 and with the inclusion of the gapped filtering 51 technique. Here, we only summarize the key components of the algorithm; see the earlier works for a more detailed explanation of the methodology. A block diagram, shown in Figure 1, depicts the major portions of the code.
@@
where $\mu$ is the chemical potential, $a_{n}$ are expansion coefficients, and $N_{chb}$ is the maximum degree of Chebyshev polynomial needed to build the filter. One requires the upper and lower spectral bounds of the Kohn-Sham Hamiltonian $\hat{H}^{KS}$ to shift-and-scale the eigenvalue spectrum into the interval $\left[-1,1\right]$; there are various schemes to obtain these bounds but we find that one of the simplest, a shifted power iteration, works well for this purpose. In the gapped-filtering method, rather than applying the cutoff of filter $\Theta$ at a specific value of $\mu$ we instead apply it over the band gap containing $\mu$, i.e. over $E_{HOMO}^{KS}\leq\mu\leq E_{LUMO}^{KS}$, so we must also map the energies $E_{HOMO}^{KS}$ and $E_{LUMO}^{KS}$ into $\left[-1,1\right]$. With these we compute the $N_{chb}$ filter coefficients $a_{n}$ as described previously 51.
@@
Subsequently, we evaluate the diagonal time-dependent self-energy matrix element, $\left\langle\phi_{k}\left|\Sigma\left(t\right)\right|\phi_{k}\right\rangle$, in two phases. In the first, we use linear-response time-dependent Hartree 48 to compute the action of the retarded polarization interaction $W^{R}$ on the occupied states. Algorithmically, we create a perturbed copy of each $\eta_{i,\ell}$ (denoted $\eta_{i,\ell}^{\lambda}$), and propagate both perturbed and unperturbed copies in time under the action of the time-dependent Hamiltonian. We accumulate the causal response function, $u^{R}\left(\mathbf{r},t\right)$, as the difference of the perturbed and unperturbed time-dependent potentials, and we time-order 55 the accumulated $u^{R}\left(\mathbf{r},t\right)$ to produce the effective polarization potential $u\left(\mathbf{r},t\right)$. Note that instead of accumulating $u^{R}\left(\mathbf{r},t\right)$ in the basis of $\left\{\eta_{i,\ell}\right\}$ (which requires keeping a copy of $u^{R}\left(\mathbf{r},t\right)$ on the full spatial grid for each $\eta_{i,\ell}$), we reduce the storage requirement by projecting $u^{R}\left(\mathbf{r},t\right)$ onto a set of $N_{\xi}$ randomly chosen functions $\left|\xi_{i,m}\right\rangle$ on short, fragmented segments of the spatial grid and accumulate $u^{R}\left(\mathbf{r},t\right)$ in the $\left\{\xi_{i,m}\right\}$ basis.
@@
where $X$ is the sROI realization of the Fock exchange operator in the basis of $\left\{\eta_{i,\ell}\right\}$ and all other quantities have been previously defined.
@@ 2.4 Utilities
The newest (3.0) version of $\mathtt{StochasticGW}$ is freely available 56 on GitHub and includes several utilities to aid researchers in preparing inputs for the code:
The $\mathtt{dft2sgw}$ utility reads and preprocesses results from a preliminary DFT calculation. This utility requires a DFT output file and a set of $\mathsf{.cube}$ files as input; $\mathtt{dft2sgw}$ prepares an input file, $\mathsf{sgwinp.txt}$, containing atomic coordinates, HOMO and LUMO energies (for gapped filtering), the spatial charge density, and a requested set of orbitals for the system of interest. $\mathtt{dft2sgw}$ also has a functionality, enabled via the $\mathtt{FFTW}$ 57 library, to up- or down-sample the orbital/density spatial grid from the preliminary DFT calculation in case a different grid for the stochastic GW step is desired. The utility currently supports Quantum ESPRESSO 25, 58, the Real-Space Multigrid ($\mathtt{RMG}$)- 59, 60 DFT code, and $\mathtt{CP2K}$ 35 (but note that $\mathtt{StochasticGW}$ does not yet include pseudopotential support for $\mathtt{CP2K}$).
$\mathtt{StochasticGW}$ also includes two utilities, $\mathtt{plotfilter.py}$ and $\mathtt{plotorbital.py}$, which use the $\mathtt{MatPlotLib}$ 61 python package to generate plots related to stochastic GW calculations:
@@ 3 Numerical Experiments
We now test our implementation of $\mathtt{StochasticGW}$ by computing QP energies of a series of non-periodic nanoclusters, $\textrm{S}\textrm{i}_{\textrm{293}}\textrm{H}_{\textrm{172}}$, $\textrm{S}\textrm{i}_{\textrm{705}}\textrm{H}_{\textrm{300}}$, $\textrm{S}\textrm{i}_{\textrm{5031}}\textrm{H}_{\textrm{1172}}$, $\textrm{S}\textrm{i}_{\textrm{7745}}\textrm{H}_{\textrm{1572}}$, $\textrm{S}\textrm{i}_{\textrm{8381}}\textrm{H}_{\textrm{1620}}$. We constructed each cluster from a uniformly expanded silicon superlattice of size $15\times 15\times 15$ using the experimental unit cell parameter for silicon ($a=b=c=10.26$ Bohr) corresponding to the diamond cubic structure with an eight-atom unit cell 62. We then shifted the coordinate origin to the geometric center of the superlattice and applied a spherical truncation, retaining only Si atoms within 20-70 Bohr of the origin. The truncated cluster is passivated with hydrogen atoms to saturate dangling bonds, and the resulting structure is relaxed to its equilibrium geometry using the MMFF94 force field 63 as implemented in Open Babel software 64. Figure 2 depicts the largest cluster, $\textrm{S}\textrm{i}_{\textrm{8381}}\textrm{H}_{\textrm{1620}}$.
@@
We performed the initial periodic DFT calculations to generate the orbitals and charge densities using the $\mathtt{RMG}$ 59, 60 DFT code. The DFT Hamiltonian uses the GGA PBE exchange-correlation functional with Troullier-Martins 65 norm-conserving pseudo-potentials. For each system, we performed the calculation on the Gamma k-point in a primitive cubic cell with isotropic sampling. Each cluster is separated from its periodic image by a vacuum layer of 11-17 bohr. The initial DFT step provides the energy estimates $E_{HOMO}^{KS}$ and $E_{LUMO}^{KS}$ used to define the gapped filter for the GW step. Cell and grid parameters, along with HOMO and LUMO energies, are listed in Table 1.
@@
From Figure 5, one can see that the portions of the code that have been ported to GPUs collectively account for $\sim$38 % of the total runtime. The remaining, unported portion of the code includes I/O operations, preparation of the grid representation of the Hamiltonian, constructing the $\left\{\zeta\right\}$ and $\left\{\xi\right\}$ orbitals prior to filtering, solving the linear system to generate the filter coefficients 51, and collecting and post-processing the samples to produce the final QP energies.
@@ 4 Conclusion
In this work, we describe a new implementation of the $\mathtt{StochasticGW}$ code. Our code utilizes the stochastic Resolution of the Identity (sROI) technique, which allows one to decouple the main steps of the GW method into independent, statistical operations that can be performed massively in parallel. In deterministic GW methods, the cost is dominated by computing matrix elements over indices representing the occupied and unoccupied orbitals. In contrast, in the stochastic method the cost depends on operations over the full spatial grid and accumulating a sufficient number of Monte Carlo samples to achieve a desired statistical accuracy. Therefore, compared to deterministic GW, the cost of stochastic GW grows much more slowly with respect to the size of the molecule or material system of interest.

### v2

@@ 1 Introduction
For systems containing thousands of atoms or more, one can evaluate the self-energy operator using a stochastic GW formulation at greatly reduced cost, as detailed by our previous works 48, 49, 50, 51, 52, 53, 40. Here, we briefly summarize the main features of the method. First, we evaluate the self-energy operator in the time domain to exploit the direct product computation of $\varSigma\left(t\right)$ from Green’s function $G$ and screened Coulomb potential $W$; we Fourier transform the resulting $\Sigma\left(t\right)$ to $\Sigma\left(\omega\right)$ only in the final stage of the calculation. Second, we invoke the stochastic Resolution of Identity (sRoI) 54, 48, 50 and define random orbital functions to use as bases for evaluating the Green’s function $G$ and effective polarization $W$. We then compute the expectation values of these operators using real time propagation and accumulate statistical averages over products of random samples. This is the main ingredient of stochastic GW and it has the advantage of allowing one to decouple the spatial- and time- dependence in the six-dimensional integrals needed to evaluate $\Sigma\left(t\right)$ 48. As a result, instead of requiring the full space of occupied and unoccupied orbitals and energies $\left\{\phi_{k},\varepsilon_{k}\right\}$ (which typically number in the tens of thousands or for a thousand-atom molecule), we, in effect, evaluate $G$ and $W$ using compact sets of stochastic linear combinations of the occupied or unoccupied orbital space. An additional benefit of sROI is that computations in the stochastic bases can be done independently, enabling the critical path of the calculation to be made embarrassingly parallelizable. Third, we incorporate sparse stochastic compression 50, 55, 56, 57, 58, 59, 60 in our stochastic Time-Dependent Hartree propagation algorithm 48, 61 for evaluating $W$. This enables computing components of $W$ over a collection of randomly-chosen short segments without needing a full spatial grid, reducing storage costs. Finally, instead of projecting each stochastic sample onto the full set of occupied orbitals $\left\{\phi_{k}^{KS}\right\}_{occ}$ from the preliminary DFT calculation (incurring substantial I/O and computational costs), we filter these samples to generate occupied stochastic orbitals. We construct a filter from a Chebyshev polynomial expansion of the Kohn-Sham Hamiltonian 62. While the filtering approach has the disadvantage of requiring many terms to produce a sharp cutoff at the Fermi energy, we recently found that the expansion length of the filter can be greatly decreased 53 by relaxing the expansion to have zero weights inside the band gap (where no states are present); this, in turn, reduces the number of matrix-vector products needed to prepare the occupied stochastic orbitals.
@@ 2.1 Algorithmic Overview
Our stochastic GW implementation is similar to that described previously 50 and with the inclusion of the gapped filtering 53 technique. Here, we only summarize the key components of the algorithm; see the earlier works for a more detailed explanation of the methodology. A block diagram, shown in Figure 1, depicts the major portions of the code.
@@
where $\mu$ is the chemical potential, $a_{n}$ are expansion coefficients, and $N_{chb}$ is the maximum degree of Chebyshev polynomial needed to build the filter. One requires the upper and lower spectral bounds of the Kohn-Sham Hamiltonian $\hat{H}^{KS}$ to shift-and-scale the eigenvalue spectrum into the interval $\left[-1,1\right]$; there are various schemes to obtain these bounds but we find that one of the simplest, a shifted power iteration, works well for this purpose. In the gapped-filtering method, rather than applying the cutoff of filter $\Theta$ at a specific value of $\mu$ we instead apply it over the band gap containing $\mu$, i.e. over $E_{HOMO}^{KS}\leq\mu\leq E_{LUMO}^{KS}$, so we must also map the energies $E_{HOMO}^{KS}$ and $E_{LUMO}^{KS}$ into $\left[-1,1\right]$. With these we compute the $N_{chb}$ filter coefficients $a_{n}$ as described previously 53.
@@
Subsequently, we evaluate the diagonal time-dependent self-energy matrix element, $\left\langle\phi_{k}\left|\Sigma\left(t\right)\right|\phi_{k}\right\rangle$, in two phases. In the first, we use linear-response time-dependent Hartree 48 to compute the action of the retarded polarization interaction $W^{R}$ on the occupied states. Algorithmically, we create a perturbed copy of each $\eta_{i,\ell}$ (denoted $\eta_{i,\ell}^{\lambda}$), and propagate both perturbed and unperturbed copies in time under the action of the time-dependent Hamiltonian. We accumulate the causal response function, $u^{R}\left(\mathbf{r},t\right)$, as the difference of the perturbed and unperturbed time-dependent potentials, and we time-order 63 the accumulated $u^{R}\left(\mathbf{r},t\right)$ to produce the effective polarization potential $u\left(\mathbf{r},t\right)$. Note that instead of accumulating $u^{R}\left(\mathbf{r},t\right)$ in the basis of $\left\{\eta_{i,\ell}\right\}$ (which requires keeping a copy of $u^{R}\left(\mathbf{r},t\right)$ on the full spatial grid for each $\eta_{i,\ell}$), we reduce the storage requirement by projecting $u^{R}\left(\mathbf{r},t\right)$ onto a set of $N_{\xi}$ randomly chosen functions $\left|\xi_{i,m}\right\rangle$ on short, fragmented segments of the spatial grid and accumulate $u^{R}\left(\mathbf{r},t\right)$ in the $\left\{\xi_{i,m}\right\}$ basis.
@@
where $X$ is the sROI 48, 50 realization of the Fock exchange operator in the basis of $\left\{\eta_{i,\ell}\right\}$ and all other quantities have been previously defined.
@@ 2.4 Utilities
The newest (3.0) version of $\mathtt{StochasticGW}$ is freely available 64 on GitHub and includes several utilities to aid researchers in preparing inputs for the code:
The $\mathtt{dft2sgw}$ utility reads and preprocesses results from a preliminary DFT calculation. This utility requires a DFT output file and a set of $\mathsf{.cube}$ files as input; $\mathtt{dft2sgw}$ prepares an input file, $\mathsf{sgwinp.txt}$, containing atomic coordinates, HOMO and LUMO energies (for gapped filtering), the spatial charge density, and a requested set of orbitals for the system of interest. $\mathtt{dft2sgw}$ also has a functionality, enabled via the $\mathtt{FFTW}$ 65 library, to up- or down-sample the orbital/density spatial grid from the preliminary DFT calculation in case a different grid for the stochastic GW step is desired. The utility currently supports Quantum ESPRESSO 25, 66, the Real-Space Multigrid ($\mathtt{RMG}$)- 67, 68 DFT code, and $\mathtt{CP2K}$ 35 (but note that $\mathtt{StochasticGW}$ does not yet include pseudopotential support for $\mathtt{CP2K}$).
$\mathtt{StochasticGW}$ also includes two utilities, $\mathtt{plotfilter.py}$ and $\mathtt{plotorbital.py}$, which use the $\mathtt{MatPlotLib}$ 69 python package to generate plots related to stochastic GW calculations:
@@ 3 Numerical Experiments
We now test our implementation of $\mathtt{StochasticGW}$ by computing QP energies of a series of non-periodic nanoclusters, $\textrm{S}\textrm{i}_{\textrm{293}}\textrm{H}_{\textrm{172}}$, $\textrm{S}\textrm{i}_{\textrm{705}}\textrm{H}_{\textrm{300}}$, $\textrm{S}\textrm{i}_{\textrm{5031}}\textrm{H}_{\textrm{1172}}$, $\textrm{S}\textrm{i}_{\textrm{7745}}\textrm{H}_{\textrm{1572}}$, $\textrm{S}\textrm{i}_{\textrm{8381}}\textrm{H}_{\textrm{1620}}$. We constructed each cluster from a uniformly expanded silicon superlattice of size $15\times 15\times 15$ using the experimental unit cell parameter for silicon ($a=b=c=10.26$ Bohr) corresponding to the diamond cubic structure with an eight-atom unit cell 70. We then shifted the coordinate origin to the geometric center of the superlattice and applied a spherical truncation, retaining only Si atoms within 20-70 Bohr of the origin. The truncated cluster is passivated with hydrogen atoms to saturate dangling bonds, and the resulting structure is relaxed to its equilibrium geometry using the MMFF94 force field 71 as implemented in Open Babel software 72. Figure 2 depicts the largest cluster, $\textrm{S}\textrm{i}_{\textrm{8381}}\textrm{H}_{\textrm{1620}}$.
@@
We performed the initial periodic DFT calculations to generate the orbitals and charge densities using the $\mathtt{RMG}$ 67, 68 DFT code. The DFT Hamiltonian uses the GGA PBE exchange-correlation functional with Troullier-Martins 73 norm-conserving pseudo-potentials. For each system, we performed the calculation on the Gamma k-point in a primitive cubic cell with isotropic sampling. Each cluster is separated from its periodic image by a vacuum layer of 11-17 bohr. The initial DFT step provides the energy estimates $E_{HOMO}^{KS}$ and $E_{LUMO}^{KS}$ used to define the gapped filter for the GW step. Cell and grid parameters, along with HOMO and LUMO energies, are listed in Table 1.
@@
From Figure 5, one can see that the portions of the code that have been ported to GPUs collectively account for $\sim$38 % of the total runtime. The remaining, unported portion of the code includes I/O operations, preparation of the grid representation of the Hamiltonian, constructing the $\left\{\zeta\right\}$ and $\left\{\xi\right\}$ orbitals prior to filtering, solving the linear system to generate the filter coefficients 53, and collecting and post-processing the samples to produce the final QP energies.
@@ 4 Conclusion
In this work, we describe a new implementation of the $\mathtt{StochasticGW}$ code. Our code utilizes the stochastic Resolution of the Identity (sROI) 48, 50 technique, which allows one to decouple the main steps of the GW method into independent, statistical operations that can be performed massively in parallel. In deterministic GW methods, the cost is dominated by computing matrix elements over indices representing the occupied and unoccupied orbitals. In contrast, in the stochastic method the cost depends on operations over the full spatial grid and accumulating a sufficient number of Monte Carlo samples to achieve a desired statistical accuracy. Therefore, compared to deterministic GW, the cost of stochastic GW grows much more slowly with respect to the size of the molecule or material system of interest.